OpenStudy (anonymous):

Evaluate the integral F(x,y) = y^2 i - xy j and C is part of the circle x^2+y^2 that starts at (1,0) and ends at (0,1)

OpenStudy (anonymous):

\[\int\limits_{0}^{1}(\int\limits_{0}^{1}(x^2 + y^2 dx))dy = \int\limits_{0}^{1}(1\div 3 + y^2 dy) = 2 \div 3\]